A) μ 0.90 B).-I C) H 1.10 D)μ-1.15 Use Binomial Distribution to solve problems 4-5 4. In a study of brand recognition, 70% of consumers recognized Coke. Find the probability that among 5 randomly selected consumers, exactly 3 will recognize Coke. (Use binomial distribution with parameters n 5 and p 0.70). Show the result of calculations and circle one of the multiple choice answers. (6 points) A) 0.6174 B) 0.2161 C) 0.3087 D) 1.0290 5. Ifa person makes 4 shots (n 4) and in each of them he has 75% chance to hit ata (p 3/4), what is the probability that he will hit the target at least two times (P(x 2 Show the result of calculations and circle one of the multiple choice answers.7 23
Answer #1

P(X=x)=n_{C_{x}}(1-p)^{n-x}(p)^{x}

4)P(X = 3) = 5G (1-0.70)-(0.70)

  10(0.30)2(0.70)

0.3087 C

5) P(Xgeq 2)=P(X=2)+P(X=3)+P(X=4)

  4(,-(0.25)4-2(0.75 )2+4C.(0.25 )4-3 (0,75). 4c. (O.25 )4-4(0.75)4

6(0.25)2(0.75)2 4(0.25) (0.75)51(0.25) (0.75)4

  0.9492